OpenStudy (anonymous):

how to take the integral of square root (1 + 2t + t^2) dt?

OpenStudy (anonymous):

$$\int\sqrt{1+2t+t^2}dt=\int\sqrt{(1+t)^2}dt=\int(1+t)dt=t+\frac12t^2+C$$